### v2

# Orders in Chaos: Enhancing Large-Scale MoE LLM Serving with Data Movement Forecasting

Abstract
Large-scale Mixture of Experts (MoE) Large Language Models (LLMs) have recently become the frontier open weight models, achieving remarkable model capability similar to proprietary ones. But their random expert selection mechanism introduces significant data movement overhead that becomes the dominant bottleneck in multi-unit LLM serving systems.
To understand the patterns underlying this data movement, we conduct comprehensive data-movement-centric profiling across four state-of-the-art large-scale MoE models released in 2025 (200B-1000B) using over 24,000 requests spanning diverse workloads. We perform systematic analysis from both temporal and spatial perspectives and distill six key insights to guide the design of diverse future serving systems. With our insights, we then demonstrate how to improve wafer-scale GPUs as a case study, and show that minor architectural modifications leveraging the insights achieve substantial performance gains, delivering 5.3× and 3.1× average speedups on DeepSeek V3 and Qwen3 respectively. Our work presents the first comprehensive data-centric analysis of large-scale MoE models and a concrete design study using the learned lessons, with profiling traces and simulation framework already open-sourced with 1k downloads. Our traces and results are publicly available at https://huggingface.co/datasets/core12345/MoE_expert_selection_trace.
[…]
I Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse domains, including programming assistance [llm_programming1, llm_programming2], translation [gpt_llm_trans1, lu2024llm_trans2], and chatbots [llm_chatbot1, llm_chatbot2]. Since the beginning of 2025, large-scale Mixture of Experts (MoE) LLMs (200B+ model with 100+ experts) have become the leading models for frontier LLMs [chiang2024chatbot] and the most widely used open weight models.
Unlike dense LLMs that activate all model weights uniformly, MoE models dynamically route each token to only a subset of experts, introducing substantial data movement overhead. Such overhead already exceeds 50% of execution time for small models (e.g., Mixtral 8x7B) on modest systems (2-4 GPUs), and it exacerbates further with larger models, such as DeepSeek V3 with 32 experts and 15 parameters deployed on multi-node systems (32+ GPUs) [go2025moetuner, zhang2025comet]. However, as shown in Figure 1, no prior work has systematically investigated the underlying data movement patterns in large-scale MoE models. Prior studies conducted before the emergence of large-scale MoEs [fang2025klotski, tairin2025emoe, yao2024exploiting] have confined themselves to profiling much smaller models on limited hardware, typically examining only one or two LLMs while reporting simple observations without system-level insights. As parameter size and expert count surge, new patterns have emerged but remain unexplored, leaving significant optimization opportunities on the table. Thus, a comprehensive study of the underlying data movement patterns in SOTA MoE models presents a fruitful opportunity for better efficiency.
If data movement in MoE models were fully unpredictable, it would present significant challenges for deployments on multi-unit systems. From a temporal perspective, the explosive growth in expert combinations would make it impossible to prefetch, cache, or replicate experts in advance. For example, large-scale MoE models like DeepSeek V3 have combinations in expert selection. When served with host memory-offload systems, such unpredictability would result in data movement like expert migrations between GPU and host, incurring substantial overhead, as inter-unit communication becomes the primary bottleneck. From a spatial perspective, if expert selection were truly random, it would lead to severe workload imbalance across units. When queries from diverse tasks are served concurrently, the number of queries assigned to each expert would vary dramatically, creating significant workload disparities. Consequently, most units would remain idle and wait for heavily loaded units to finish, resulting in poor hardware resource utilization.
Fortunately, as we later show in the paper, MoE expert selections indeed have predictability that designers can exploit to reduce data movement. To uncover the inherent patterns in MoE models, we conduct a comprehensive data-movement-centric profiling of four state-of-the-art MoE models ranging from 235B to 1000B parameters released in 2025. As highlighted in Figure 1, we profile DeepSeek V3 [liu2024deepseek], Llama4 Maverick [llama4], Qwen3-235B [yang2025qwen3], and Kimi K2 [team2025kimi] across 24,000 requests involving varied tasks, topics, and languages, which consumes 2000 GPU hours in total. We then collect the expert selection trace of all layers and tokens in each request to create an expert selection database of over 150 GB JSON files. From these extensive traces, we conduct a comprehensive analysis to uncover data movement patterns from both temporal and spatial perspectives, making our findings system-agnostic and applicable to various serving architectures at any scale. We then distill six key insights that serve as a solid foundation to understand MoE data movement and directly inform future MoE LLM serving system design, addressing critical questions that have remained unanswered in the field, such as: Is there a correlation between previously selected experts and those selected later? Are there discernible rules underlying the observed expert selection skewness? Do different tasks tend to activate different experts? Our work represents the first systematic effort to characterize data movement patterns at the scale of up-to 1000B model across a wide range of tasks, providing actionable insights that can guide the design of next-generation MoE serving systems.
To demonstrate the broad applicability of our insights, we present a case study on future GPU architecture design. Modern GPUs have adopted multi-chiplet designs due to single-die size limitations [mi300, cpelide, blackwell] and are evolving toward wafer-scale architectures enabled by emerging on-wafer packaging technologies [he2025waferlm, xu2025wsc]. By applying our insights, we develop two key innovations for future wafer-scale GPUs: (1) A task distribution algorithm with a two-level data-placement-aware command processor architecture that balances workload across dies, and (2) a data-driven predictor with hardware-managed HBM architecture that reduces inter-die communication overhead by automatically caching critical data locally. These techniques achieve an average 4.0-6.3 speedup in MoE LLM serving throughput on wafer-scale GPUs.
Our main contributions can be summarized as follows:
We propose a comprehensive and systematic data-movement-centric profiling across four latest, large-scale MoE models released in 2025 between 235B and 1000B to uncover the data movement patterns from both temporal and spatial perspectives.
We distill six key insights for designing efficient MoE serving systems based on our profiling and analysis, providing actionable guidance that can inspire future research in MoE serving systems.
Leveraging these insights, we conduct a case study on wafer-scale GPU architecture design. We proposed two strategies that improve MoE serving performance by 4.0-6.3 with minor hardware modifications.
We collect over 70,000 expert selection traces across multiple models and datasets, totaling over 150 GB in JSON format, and have open-sourced (1k downloads) all traces with our multi-chiplet simulator to facilitate future research on MoE serving systems.
II Background
II-A LLM and MoE Model Architecture
Most state-of-the-art LLMs adopt a decoder-only transformer architecture that follows a token-by-token auto-regressive workflow [attention]. As shown in Figure 3(a), after users input queries, the serving process is divided into two stages: the prefill stage and the decode stage. During the prefill stage, all input tokens are processed simultaneously to generate the first output token. The decode stage follows immediately, where tokens are generated sequentially. The generated token from each iteration is appended to the input sequence to produce the next token in the following iteration.
The Mixture of Experts (MoE) mechanism is a state-of-the-art approach to improve LLM performance and has become prevalent among current frontier LLMs [moe-survey]. As shown in Figure 3(b), MoE-based LLMs replace the feed-forward network (FFN) layers in traditional LLMs with MoE layers. In each layer, multiple experts are deployed, and each request is routed to a small subset of the most suitable experts based on a gating mechanism. This innovation enables MoE models to scale model parameters without incurring extra inference overhead, since only a fraction of parameters are activated for each request. However, this mechanism also introduces dynamic randomness, since expert selection is unknown until gating is completed, posing new challenges for serving systems.
II-B Prior MoE Serving Systems
The MoE mechanism constitutes the primary source of data movement overhead in modern serving systems. As illustrated in Figure 2, MoE-related data movement (MoE All-to-All and MoE Weights) dominates the overhead across different DeepSeek V3 serving configurations, accounting for 60%-90% of total latency. To address this overhead, existing research on MoE serving optimization has developed numerous system-level solutions targeting different performance and cost objectives. Edge systems like MoE-Lightning [moelightning] and CoServe [suo2025coserve] employ CPU memory offloading techniques to address GPU memory capacity constraints, while server systems such as Comet [zhang2025comet] and MegaScale-Infer [zhu2025megascale-infer] leverage multi-GPU architectures and address GPU-GPU communications in MoE for higher throughput. Novel hardware architectures like Duplex [yun2024duplex] explore processing-in-memory to accelerate data movement in MoE LLMs.
However, these prior studies employ a system-centric methodology when optimizing for MoE LLMs. Namely, they inherently focus on a specific platform and the corresponding data movement patterns of MoE in such platform (e.g., CPU-GPU, multi-GPU, ML accelerators). As a result, they propose deployment-specific optimizations that may not generalize across different serving platforms, and their insights are often a slice of the overall inherent patterns in MoE LLMs.
In this work, we flip the process and adopt a model-centric strategy by conducting system-independent profiling to extract system-agnostic insights about MoE data movement patterns. These insights are therefore broadly applicable across various platforms, providing a foundation for optimization strategies that transcend specific system implementations.
III MoE Profiling and System Insights
In this section, we conduct a data-movement-centric profiling of the expert selection behavior in four state-of-the-art MoE models: Deepseek V3 (671B), Llama4-Maverick-128E (402B), Qwen3-235B (235B), and Kimi K2 (1000B). All results are averaged over more than 24,000 requests.
III-A Categorization Methodology
As in Figure 3, we categorize MoE expert selection profiling results into two categories: temporal and spatial relations.
Temporal relations capture time-dependent expert selection patterns where current choices inform future selections. These patterns enable single-unit strategies that optimize data movement within individual units through prefetching, caching, and data migration. For example, in multi-chiplet GPU systems, caching experts in local DRAM after remote fetches significantly reduces inter-unit communication. To exploit temporal predictability, we analyze expert selection at multiple time scales shown in Figure 3(a): layer-level, token-level, and prefill-decode-level patterns.
Spatial relations reflect expert selection distribution and imbalance across the system. These patterns enable multi-unit strategies that optimize system-wide expert placement and workload distribution. By identifying frequently selected experts and analyzing selection skewness, we design algorithms for workload routing and expert replication that balance traffic among units. We classify spatial relations into single-expert activation imbalance and expert-pair co-activation affinity as shown in Figure 3(b), and investigate how task types influence these patterns to inform system-level strategies.
III-B Temporal Relations
As shown in Figure 3(a), we classify the temporal relations of expert selection into three categories, arranged in order of increasing time scale. At the layer level, we examine the relationship between two adjacent model layers. At the token level, we focus on the same model layer across two adjacent tokens. At the stage level, we analyze the relationship between the prefill stage and the decode stage.
III-B1 Layer-Level
As shown in Figure 4 (a) and (b), we present heatmaps for Deepseek and Qwen illustrating expert selection relationships across adjacent layers. Each pixel in the heatmap displays the conditional probability of selecting expert in the next layer given that expert was activated in the previous layer, with bright colors indicating higher probabilities.
The heatmaps reveal clear cross-layer correlations with white dots highlighting specific expert pairs with significantly higher selection probabilities across adjacent layers. However, correlation patterns vary across layers within the same model and differ between models due to architectural variations. For instance, patterns between layers 3-4 differ from those between layers 30-31. Qwen3’s notably brighter heatmap indicates stronger cross-layer correlations than Deepseek. Beyond the white dots, there are also consistent bright vertical lines, suggesting certain experts are frequently chosen regardless of previous layer selections. These patterns indicate generally popular experts, analyzed further in Sec. III-C1.
To quantify these relationships, we analyze expert pair frequencies across multiple requests and all model layers. The cumulative percentage of the most frequently selected expert pairs is shown in Figure 4(c). The results demonstrate that the top 20% of expert pairs account for 45%, 68%, 80%, and 55% of all activations in Deepseek, Qwen, Llama4, and Kimi K2, respectively. This reveals strong, model-dependent cross-layer correlations, with Llama4 111Llama4 places dense FFN layers between MoE layers, so we analyze correlations between adjacent MoE layers ( and ). exhibiting the strongest effect and Deepseek the weakest.
III-B2 Token-Level
We examine expert selection relations for the same layer between adjacent tokens in Figure 5.
Each pixel in the heatmap displays the conditional probability of selecting expert in the next token given that expert was activated in the previous token, with bright colors indicating higher probabilities.
Similar to layer-level patterns, cross-token heatmaps exhibit white dots, bright vertical lines, and variation across layers and models, indicating correlations between adjacent tokens. However, token-level relations reveal a common pattern appearing across all models: the bright diagonal line that indicates the tendency to select the same expert across adjacent tokens. This diagonal pattern emerges predominantly in higher layers (17 and 43) but not in lower layers (1 and 3), regardless of models.
We quantify these token-level relations using the same approach as layer-level relations in subsubsection III-B1. As shown in Figure 5(d), the top 20% of expert pairs account for 40% to 80% of all selections, indicating strong token-level relations. Llama4 exhibits the strongest relation while Deepseek shows the weakest, suggesting future expert selections are most predictable in Llama4.
III-B3 Prefill-Decode-Level Correlation
Building on the layer-level and token-level relations, we observe notable similarities in expert selection patterns between prefill and decode stages. Comparing heatmaps across stages in Figure 6(a)(b)(c)(d), we find similar distributions of bright dots, indicating that expert pairs frequently selected during prefill are also likely selected during decode. This cross-stage consistency suggests expert selection in the decode stage can be predicted using prefill data.
To quantify this similarity, we compute Spearman’s ratio () across all model layers, comparing prefill and decode heatmaps. Spearman’s Ratio measures monotonic relationships between variables, ranging from (perfect negative correlation) to (perfect positive correlation). Generally, indicates strong correlation, indicates moderate correlation, and suggests weak correlation [spearman]. The results in Figure 6(e)(f) show that most layers demonstrate strong correlation, while a few show moderate correlation. This makes it possible to predict decode-stage expert selection with prefill-stage data.
III-B4 System Insights from Temporal Relation
The observed temporal relations in expert selection motivate us to design fine-grained, dynamic strategies on every single unit to reduce the communications. For example, when expert weights are read from remote memory, such as remote DRAM in multi-chiplet systems, or CXL extension memory in memory-disaggregated systems, caching, migration, and prefetching strategies can be deployed to reduce data movement.
Empirical analysis shows that experts selected more frequently during the prefill stage are also likely to be chosen during the decoding stage. Thus, the expert selection information collected during the prefill phase can serve as an important reference for expert selection prediction at the beginning of the decoding stage, when only a few tokens have been generated and reference data is limited. Furthermore, prefill information can also be used to dynamically adjust expert distribution during the decode stage. This is especially useful in modern serving systems that employ parameter disaggregation, where the prefill and decode stages may reside on different machines.
Layer-level and token-level temporal relations share similarities in definition but differ in their reuse distances, making them suitable for different memory hierarchy levels. Layer-level relation exhibits short reuse distances since consecutive MoE layers execute immediately, while the token-level relation creates longer distances as tokens are generated only after traversing all layers.
This distinction aligns naturally with the complex multi-unit MoE serving systems. For example, in multi-chiplet architectures, each die contains LLC and local DRAM, forming a two-level hierarchy. The faster LLC manages experts with short reuse distances (layer-level), while the larger DRAM handles longer distances (token-level). We can leverage layer-level relations for LLC management and token-level relations for local DRAM management.
This principle extends to other systems where layer-level relations can optimize faster memory tiers and token-level relations can enhance slower tiers: multi-GPU systems with LLC and local DRAM, CXL-based systems with local DRAM and remote CXL memory, SSD offloading systems with DRAM and flash storage, and in-flash computing systems with local and remote flash dies.
III-C Spatial Relation
As shown in Figure 3(b), we analyze spatial patterns in expert selection for both single-expert activation imbalance and expert pair co-activation affinity. For single experts, we examine statistical skewness and the factors affecting each expert’s activation. For expert pairs, we analyze co-activation properties across all two-expert combinations.
III-C1 Single Expert Activation Imbalance
We examine expert selection frequency at each layer, presenting results for layer 7 of Llama4 in Figure 7. After normalizing selection counts by the average, we observe pronounced skewness where a subset of experts is activated over 16 times more frequently than average. This workload imbalance suggests system designs should duplicate or decentralize frequently used experts.
To investigate selection patterns across different tasks, we analyze all 57 MMLU subjects spanning diverse fields, including biology, history, and math, etc [mmlu]. Figure 7(b) shows the top 10 most popular experts for each subject. Horizontal bright lines indicate certain experts are consistently activated regardless of subject, while remaining popular experts vary significantly between subjects, demonstrating both overlap and distinction in task-based expert selection.
We further examine task impact using Chinese version of MMLU in MMLU Pro [mmlu_pro] with identical questions but different languages. Figure 7(c) reveals distinctly different patterns: although 5-6 experts remain popular across subjects, only two overlap with English MMLU’s most frequently selected experts. This confirms that task characteristics, including language, significantly influence expert selection, enabling task-aware serving systems that optimize expert distribution to balance workloads and reduce data movement.
III-C2 Expert Pair Co-activation Affinify
Beyond single expert patterns, we observe spatial relations for expert pairs where certain experts are more likely to be co-activated. We present co-activation heatmaps in Figure 8(a)(b), where both axes indicate expert IDs. Each pixel represents an expert pair with values showing co-activation frequency normalized by theoretical random selection probability: , where is the number of experts.
Bright dots appear with probabilities 20-40 times higher than theoretical values, indicating strong co-activation tendencies. All heatmaps exhibit central symmetry since expert pair equals . In Deepseek’s heatmap Figure 8(a), frequently activated pairs lie between red lines forming bright squares, reflecting Deepseek’s routing restriction where tokens are routed only to adjacent nodes to reduce communication overhead. This suggests the potential of separating co-activated expert pairs to balance workload.
We quantify this relation by aggregating data across all layers in Figure 8(c). The top 10% of expert pairs account for 60-80% of total activations, indicating strong skewness in pair selection. Qwen exhibits even stronger co-activation relations than Deepseek, suggesting that expert distribution strategies separating frequently co-activated experts can significantly improve system performance.
We only analyze Deepseek and Qwen since Llama selects one expert per MoE layer, eliminating co-activation relations.
III-C3 System Insights from Spatial Relation.
Spatial Relation enables coarse-grained, static strategies to address workload imbalance across system. These strategies could be applied at system startup or during periodic redistribution (e.g., every 10 minutes) through appropriate task distribution.
Expert placement in serving systems can change dynamically due to expert migration strategies. Therefore, when allocating workload to system units, expert placement should be considered for better workload balance. Besides, the design space for task allocation could be enlarged with emerging new systems. Traditional multi-GPU systems tend to allocate experts to local GPUs to avoid cross-unit communication. However, in multi-chiplet GPUs, we can consider allocating tasks to remote dies for better workload balance as inter-unit communication becomes faster.
Expert skewness causes workload imbalance and suboptimal resource utilization. Duplicating popular experts across multiple compute units distributes load more evenly. Additionally, avoiding co-location of highly popular experts on the same unit further enhances workload balance.
Certain experts are frequently activated simultaneously, exhibiting strong co-activation patterns. Assigning these co-activated expert pairs to different compute units maximizes hardware parallelism and prevents workload concentration on specific units. However, separation also introduces cross-unit communication overhead. The effectiveness depends on system topology and batch size, requiring careful trade-off between parallelism benefits and communication costs.
Hot experts vary by task and language. English queries, for instance, activate different expert subsets than Chinese queries. Providing task metadata during serving enables proactive expert placement: when workloads are predominantly English, systems can pre-duplicate or reassign English-relevant experts, reducing communication and balancing loads. This task-to-expert mapping requires only one-time offline profiling per model and can be reused throughout deployment, making the approach practical and efficient.
IV Case Study: Future GPUs Architecture Design for MoE Models
In this section, we adopt future GPU architecture design as a use case to validate our proposed insights. We follow Insight 3 to design a task allocation algorithm and leverage the temporal relation insights (Insight 1 and Insight 2) to build a data-driven predictor. We also make slight architectural modifications to support the task allocation algorithm and the data-driven predictor.
IV-A Trend of Future GPU Architecture
GPU vendors are increasingly adopting multi-chiplet architectures to overcome single-die performance limitations. As Moore’s Law approaches its limits [lundstrom2003moore] and single-die size remains constrained by photomask dimensions (800-1,000 mm2), advanced packaging technologies like TSMC’s CoWoS [hu2023cowos], Samsung’s X-Cube [xcube], and Intel’s EMIB [emib] enable multiple chiplets within a single package. Leading vendors have adopted such designs: AMD’s MI300 [mi300] integrates eight compute chiplets, NVIDIA’s Blackwell features two chiplets [blackwell], and upcoming Rubin expects four [rubin].
This trend is evolving toward wafer-scale systems [cowos2]. TSMC’s System-on-Wafer (SoW) technology accommodates up to 24 compute dies and 96 HBM dies on a single wafer, exceeding 200,000 mm2 [sow-x]. As shown in Figure 9(a), a typical wafer-scale multi-chiplet GPU consists of multiple units, each containing a GPU die and several HBM dies interconnected in a mesh topology. Such systems contain over 10 TB of HBM and PFLOPS-level computing power, supporting extremely large models and batch sizes.
The TSMC SoW technology shown in Figure 9(b) connects each GPU die to local HBM dies through local-silicon interconnects (LSI) [lsi], with adjacent GPU dies communicating via LSI vertically or XSR SerDes links horizontally. Although LSI and SerDes both provide terabit-level bandwidth, inter-GPU-die communication remains the primary bottleneck. Remote data access requires multiple hops across die-to-die links, resulting in high latency. Simultaneous remote HBM access by multiple dies creates bandwidth contention and traffic congestion, further degrading performance.
IV-B Challenges of serving MoE with future GPUs
Unlike current multi-GPU systems, wafer-scale GPUs can fit entire MoE models on a single chip and support batch sizes over 10,000. However, current GPU architectures introduce two key limitations for such large-scale chips.
Simplistic Task Allocation. Current GPUs integrate a CPU in their SoC to serve as a command processor and allocate tasks to all SMs. However, the traditional command processors treat all SMs equally, ignoring their physical locations and data placement [cp1, cp2]. This oblivious task-to-SM assignment generates excessive D2D traffic and ignores MoE expert selection skewness, leading to poor utilization when most dies remain idle while others become overloaded.
Inadequate Local HBM Management. Current GPUs treat all HBM dies as uniform memory space, but wafer-scale GPUs connect each compute die directly to local HBM, where access is significantly faster than a remote HBM. Frequently accessed experts in remote HBM could be cached locally to minimize D2D traffic, but current GPUs do not distinguish between local and remote HBM and therefore generate unnecessary traffic.
IV-C Motivation and Insights
To address these challenges, we propose two strategies with architectural support. First, based on Insight 3 that identifies the need for expert-placement-aware task distribution, we propose an intelligent task allocation algorithm with a multi-level, data-placement-aware command processor architecture. This approach considers expert placement and selection skewness across dies, enabling dynamic task allocation that minimizes D2D traffic while balancing workload.
Second, leveraging Insight 1 and Insight 2 that reveal the predictability behind expert selection across different timescales, we introduce a data-driven predictor with hardware-managed HBM architecture. Local HBM caches frequently accessed experts from remote dies, while a lightweight predictor analyzes selection patterns to estimate future needs, caching predicted experts locally to reduce D2D traffic.
These enhancements work independently, each delivering significant improvements with minor hardware modifications. Combined, they provide up to 20% additional benefit over individual deployment. We implement both strategies through architectural changes to maintain consistency with existing GPU programming models, which treat GPUs as single-die devices without exposing die-level control. If future programming models provide more fine-grained features, these strategies could alternatively be implemented at the system level without architectural changes.
IV-D Architecture Design
IV-D1 Architecture Overview
To support our proposed predictor and task allocation algorithm, we implement two architectural modifications with minimal overhead, as illustrated in Figure 9(a). These changes, highlighted in orange, consist of an enhanced Command Processor structure and an extended D2D controller design.
First, we redesign the Command Processor (CP) with a two-level hierarchical structure: a Global CP at the wafer level and Local CPs within each die. The Global CP maintains system-wide expert selection and placement information for intelligent resource management. Second, we extend the D2D controller with an Address Translation Unit (ATU) and a Prediction Unit (PDU). The ATU translates remote HBM addresses to local addresses when remote data is duplicated, while the PDU identifies important remote data requiring duplication. These enable autonomous caching of remote data in local HBM and intelligent redirection of data requests, reducing inter-die communication overhead.
IV-D2 Key Data Structures
There are two key data structures: Global CP data and the PDU prediction table.
Global CP Data Structures: As shown in Figure 9(c), the Global CP maintains two structures. The expert distribution table stores each expert’s initial die ID and distribution status as an -bit binary code, where each bit indicates expert presence on the corresponding die. The cross-token heatmap records expert activation patterns over time, providing historical data for prediction generation.
PDU Prediction Table: Each PDU stores a prediction table with two key fields per expert: the cp_en bit indicating whether the expert should be cached locally (calculated by Global CP and transferred before computation), and the is_local bit tracking whether the expert is already cached (maintained by PDU itself).
IV-D3 Workflow During Kernel Launch
When a new kernel launches ( ), the Global CP implements our task allocation algorithm to split the MoE kernel into sub-kernels for each die and executes the predictor algorithm to generate duplication guidance (cp_en field is PDU). The Global CP then sends sub-kernels and prediction information to local CPs ( ). After receiving the guidance information from Global CP, each local CP allocates tasks to its SMs ( ) and configures the prediction table in the D2D controller for local HBM management ( ). After computation, the local CP collects expert duplication statistics and sends them to the Global CP to update expert distribution information.
IV-D4 Dataflow for Remote Data Access
We integrate ATU and PDU into the D2D controller to support hardware-managed HBM by modifying the remote data access flow. With these two units, our architecture automatically duplicates important remote data in local HBM, with green lines representing non-duplicated data reads and blue lines representing duplicated data reads, as shown in Figure 9(a).
Remote Data Read (Non-duplicated): When an SM reads remote data not in local HBM ( ), the D2D controller routes the request conventionally ( ). Upon return, the PDU checks the Prediction Table to make a duplication decision and sends data to the SM regardless of the decision ( ). If duplication is required, the PDU writes to LLC and local HBM ( , ), updates the address into ATU, and sets the is_local bit in PDU’s Prediction Table to 1.
Local Data Read (Duplicated): When an SM reads remote data already cached in local HBM ( ), the ATU translates the remote address to a local address and redirects the request to LLC ( ). The LLC and memory controller then retrieve data and send it back to the SM( , ).
IV-D5 Algorithm Design
This subsection presents our task allocation algorithm and data-driven predictor, both implemented by the global CP.
Task Allocation Algorithm: Since accurate task distribution is NP-hard, we propose two heuristic mechanisms: a candidate mechanism that reduces the number of dies to consider and a block-granularity distribution mechanism that searches for approximate solutions among candidates.
This algorithm splits MoE kernel computation into sub-tasks for individual dies based on expert selection and distribution information. As shown in algorithm 1, the input contains the number of requests belonging to each expert, while provides dynamic expert distribution information from the Expert Distribution Table in Figure 9(c), indicating where each expert is stored.
The algorithm iterates through all experts to generate allocation plans. For each expert, it creates a candidate die list including dies storing expert weights and their adjacent dies (shown as green and red in Figure 10(a)). We sort candidates by workload and limit the list to dies, determined by the expert’s request count (line 3-5). Requests are distributed to candidate dies in blocks of size 50 to balance efficiency and accuracy (line 6-11). For each block, the algorithm selects the optimal die using our cost model, which considers DRAM access, computation, and die-to-die communication. Finally, we merge blocks allocated to the same die to generate the final allocation plan (line 12).
Data-Driven Predictor: Our predictor algorithm, implemented by the global CP, uses current MoE kernel expert selection information to predict popular experts for the next token on each die. This prediction result is transferred to local CPs and configured in each die’s PDU to guide hardware-managed local HBM. As shown by the red shadow in Figure 10(b), we first identify corresponding rows from the heatmap based on current expert selection ( ), then select the top experts from each row ( ) and identify corresponding experts for the next token as prediction results, denoted by the green shadow ( ). In this example, a die computes experts 1 and 4 during the current MoE kernel and we predict experts 2, 4, and 6 will be used next. Since this die only reads experts 1 and 4 currently, we duplicate only expert 4 in its local DRAM.
V Evaluation
V-A Experiment Setup
Methodology: We conduct experiments using trace-driven simulation on a validated simulator. Expert selection traces are collected by deploying SGLang [zheng2024sglang] on an 8×H100 DGX server and an 8×H200 AWS instance.
We developed a custom multi-chiplet GPU simulator in Python since existing tools are inadequate for our needs. Cycle-accurate simulators like Gem5 [gem5], gpgpusim [gpgpusim], and mgpusim [mgpusim] accurately model single GPUs but are too slow for large-scale systems with 20+ GPUs serving batch sizes exceeding 15,000. Event-driven ASTRA-sim [astra] supports multi-GPU systems but lacks MoE serving support and detailed single-GPU microarchitecture modeling. To ensure accuracy, we validated our simulator against real data from an 8×H100 DGX server, as shown in Sec. V-B.
Our simulator models multi-chiplet GPU components, including LLC, HBM, compute units, and D2D links across all dies. A central resource manager simulates contention and congestion by dividing each expert execution into multiple events and managing simultaneous execution of thousands of events. The simulator supports configurable die numbers, placement, and connectivity. To balance accuracy and efficiency, we simulate MoE models at expert slice granularity, with each expert comprising two slices.
| X-die | Y-die |
DRAM |
| Dojo | 5 | 5 | 2 TB/s | 1.5 TB/s | 80GB | 1000 TFLOPS | ||||||
| TSMC-SoW | 3 | 8 | 2 TB/s | 1.5 TB/s | 80GB | 1000 TFLOPS | ||||||
| Dojo-Enhanced | 5 | 5 | 8 TB/s | 2 TB/s | 180GB | 4500 TFLOPS | ||||||
| Other Params |
Metric: We measure pure MoE layer performance during the decode stage as modern LLM serving systems show a trend toward fine-granularity disaggregation. Traditional LLMs benefit from separating prefill and decode stages across different machines, as demonstrated by DistServe [zhong2024distserve] and subsequent works [loongserve, qin2025mooncake]. For MoE models, this disaggregation extends further. MegaScale-Infer [zhu2025megascale-infer] separates attention and MoE operations onto different machines for optimal batch sizes. Following this trend, we focus on optimizing MoE operations during the decode stage.
Hardware Configuration: We evaluate two multi-chiplet topologies: Tesla Dojo [dojo1, dojo2] and the TSMC SoW roadmap [tsmc-arch]. As summarized in Table I, Dojo uses a 55 2D mesh, while TSMC SoW adopts an 83 2D mesh. These choices reflect a deployed system (Dojo) and near-future industry support (TSMC SoW).
For both the Dojo and TSMC SoW configurations, each chiplet is A100-like, providing 1,000 TFLOPS FP8 compute, 80GB HBM, 2TB/s local HBM bandwidth, and 1.5 TB/s inter-die bandwidth to adjacent chiplets. We also include an extended experiment in Sec. V-F with a Dojo-Enhanced configuration, where each die is B300-like to reflect an anticipated hardware performance trend in the future. We reserve 10% of DRAM for system and hardware management.
Baseline Configurations: We compare our approach against the simple strategy currently used by GPU. The Base configuration assigns equal numbers of experts to each die but ignores expert placement when allocating tasks. We implement three variants: Allo Only uses solely our task allocation strategy; Pred Only includes only the data-driven predictor; and Allo+Pred combines both techniques. These configurations evaluate the individual and combined effects of our proposed methods.
Models and Workloads: We conduct evaluations with real traces collected from Qwen3 and Deepseek V3. The traces are gathered from diverse datasets, including MMLU [mmlu], MMLU Pro [mmlu_pro], ChineseSimpleQA [he2024chinese], and LiveCodeBench [jain2024livecodebench], comprising over 24,000 requests per model.
V-B Validation of Simulator
We validate our simulator using real measurements from an 8×H100 DGX server. We evaluate both single-GPU execution and two-GPU peer-to-peer (P2P) communication. For single-GPU execution, we benchmark GEMM operations for experts in the MoE layer across varying batch sizes for both DeepSeek and Qwen. For P2P communication, we measure data migration between two GPUs over payload sizes ranging from 4 KB to 4 GB. To ensure simulation fidelity, we calibrate key parameters to fit the measured data. As shown in Figure 12, the simulator’s error remains within 5% for all test cases.
V-C Throughput
We evaluate MoE decode stage throughput in Figure 11, with results normalized to the baseline configuration.
Comparison across models: Our Allo+Pred strategy achieves 5.3 throughput improvement on Deepseek and 3.1 on Qwen. Deepseek shows greater improvement due to its larger number of experts (256 vs. 128) and more complex expert selection patterns. Without our expert placement-aware allocation, serving Deepseek introduces significantly more inter-unit communication than Qwen. Since our task allocation strategy places computation close to data, it yields more substantial improvements on Deepseek.
Comparison across chiplet architectures: Our Allo+Pred strategy shows 3.7 improvement on Dojo and 4.7 on TSMC, despite similar die counts (25 vs. 24). The improvement on TSMC is more substantial due to its rectangular layout, where dies are farther apart, introducing more inter-unit communication without strategic task allocation.
V-D Hop Reduction
We report hop counts in Figure 11 to show the reduction in inter-unit communication. Hop count is the sum of Manhattan distances for all cross-unit communications. Higher hop counts indicate frequent cross-die data movement. We normalize results to baseline and report hop reduction ratios, where a ratio of 10 means hop count is reduced to 1/10.
Pred Only reduces hop counts by 3.7-4.4, aligning with performance results in Figure 11. This indicates cross-unit communication is the primary bottleneck in baseline, and reducing hop counts proportionally improves performance.
Allo Only reduces hop counts by 7.2-10.8, exceeding the performance improvement range of 17-661. This shows that with our allocation algorithm, inter-unit communication is no longer the sole bottleneck. While reducing hop counts still improves performance, the improvement is not proportional.
Allo+Pred reduces hop counts by over 210-944 compared to baseline and achieves 10 improvement over Allo Only. However, performance improvement is only 2.5-6.5 over baseline, with just 1.2 average improvement over Allo Only. This demonstrates that hop count is no longer a performance bottleneck. Our task allocation successfully places computation near relevant data. Most tasks are distributed to local dies holding related experts, with only extremely popular experts requiring remote allocation. This leads to minimal D2D traffic and shifts the bottleneck to workload distribution.
V-E DRAM Access Breakdown
We provide a breakdown of DRAM access patterns in Figure 14 to show how our strategies reduce inter-unit communication. We categorize DRAM access into three types: reads from local dies, reads from remote dies, and writes to local dies, where writes to local dies only occur when we duplicate a remote expert locally. Most reads in the baseline are from remote dies, resulting in high inter-unit traffic and poor performance. With our strategies (Pred Only, Allo Only, and Allo+Pred), most remote DRAM reads are converted to local DRAM reads, significantly reducing traffic. Compared with Pred Only, Allo+Pred achieves fewer remote reads by allocating most tasks to local dies, with only extremely popular experts requiring computation across multiple dies. Compared with Allo Only, Allo+Pred further reduces remote reads by caching popular experts in local HBM.
V-F Comparison with Host CPU-Based Implementation
Our task allocation algorithm runs on a new GPU command processor, but it could, in principle, be executed on the host CPU with a higher overhead. As shown in Figure 14, we evaluate both the Dojo and Dojo-Enhanced configurations. In Dojo, the overhead of host-CPU allocation is 5.2%–6.4% for DeepSeek V3 and 11.1%–14.2% for Qwen3. In Dojo-Enhanced, the overhead rises to 19.3%–23.8% for DeepSeek V3 and 42.0%–51.6% for Qwen3.
DeepSeek vs Qwen: Qwen3 incurs higher overhead than DeepSeek V3 due to CPU–GPU data transfers over PCIe, which occur once per MoE layer. The CPU needs the Expert Distribution Table from the GPU to run the allocator, and the allocation results must be sent back to the GPU before kernel execution. Qwen3 has (i) more MoE layers (94 vs. 58), increasing transfer frequency, and (ii) smaller per-layer compute, which amplifies the relative cost of transfers.
Dojo vs Dojo-Enhanced: Dojo-Enhanced shows over 3.7× higher overhead than Dojo because its GPU dies are significantly faster, making fixed PCIe transfer costs dominate more. As GPU performance outpaces interconnect bandwidth, implementing the allocator in the GPU command processor becomes increasingly necessary to sustain performance.
V-G Discussion
This wafer-scale GPU architecture serves as a case study demonstrating the practical applicability of our proposed insights, which constitute the paper’s primary contribution. Our design implements only a subset of the proposed insights to validate their effectiveness. We follow Insight 3 for task allocation and leverage temporal relation insights (Insight 1 and part of Insight 2) to build a data-driven predictor. However, the remaining insights offer significant potential for further exploitation in wafer-scale architectures. For instance, Insight 2 could inform LLC prefetch strategies and enhance task allocation algorithms that consider LLC data placement. Spatial relation insights (Insight 4 and Insight 5) could optimize initial expert placement, which could be further refined by incorporating task information from Insight 6. Importantly, our insights extend far beyond this case study to benefit diverse future systems, including multi-GPU clusters [zhu2025megascale-infer, zhang2025comet], CXL-based or CPU-based memory disaggregation [moelightning, fang2025klotski], Flash-based multi-hierarchy systems [cambricon-llm, sheng2023flexgen], and other emerging architectures and systems.
VI Related Work
MoE model behavior studies:
Several MoE model tech reports [muennighoff2024olmoe, jiang2024mixtral, dai2024deepseekmoe, sun2024hunyuan] provide the MoE routing patterns as part of their evaluation.
For example, the Mixtral report [jiang2024mixtral] shows the temporal locality of expert selection by reporting the percentage of repetitive assignment.
The OLMoE report [muennighoff2024olmoe] shows the co-activation pattern and domain specialization among the experts.
A blog post from SGLang [sglang-deepseek-blogpost] shows expert distribution statistics for the DeepSeekV3 model and the inherent imbalance in expert selection and similarity between prefill and decode.
None of these studies provides a comprehensive profiling across multiple large-scale (200B) MoE models, nor do they present a data-movement-centric methodology to highlight the opportunities like this paper.
Data movement optimization for MoE LLM inference
Various prior work [hwang2024pre, li2023accelerating, zhai2023smartmoe, gupta2024lynx, kamahori2025fiddler, eliseev2023fast, du2024sida, li2024optimizing, skliar2024mixture, moelightning, yun2024duplex, zhang2025comet, hwang2023tutel] have focused on improving the efficiency of MoE LLM inference by reducing data movement.
For example, Lina [li2023accelerating] exploits the skewness in expert selection to dynamically schedule resources during inference and balance traffic across multiple GPUs.
LYNX [gupta2024lynx] dynamically reduces active experts while preserving model accuracy.
Pre-gate MoE [hwang2024pre] uses a pre-gating function to alleviate the dynamic nature of expert selection.
Sida [du2024sida] builds an offline hash function to predict expert usage to reduce data movement between CPU and GPU.
MoE-Lightning [moelightning] leverages a CPU-GPU pipeline and paged weights to improve resource utilization.
Eliseev and Mazur [eliseev2023fast] exploit expert locality and leverage LRU caching to manage GPU and CPU memory.
This work aims to reduce data movement in MoE LLM inference of today’s systems, and our case study shows that data-driven profiling across multiple models provides a solid base to optimize not only current but future systems of any scale.
Wafer-Scale and Chiplet Architectures
As single chip scaling slows down, packaging technologies such as wafer-scale and chiplet packaging have become a promising path to further improve compute efficiency for future systems.
Recent work in this area either focuses on the interconnect design [rashidi2025fred, yang2025pd, yu2025cramming, li2024lucie, chen2024waferscale, feng2023heterogeneous, feng2023scalable] of wafer-scale and chiplet systems, or
general data placement for specific algorithms [he2025waferlm, tan2021nn] and applications [shao2019simba, xu2025wsc, odema2024scar].
In contrast, our work is the first to study the implications of serving large-scale MoE LLMs in a wafer-scale GPU and propose data-movement-centric HW/SW co-design optimizations.
VII Conclusion
In this work, unlike previous MoE serving works that adopt system-centric approaches and develop deployment-specific strategies, we study MoE model serving from a model-focused perspective We conducted comprehensive data-movement-centric profiling of the latest large-scale MoE models (200B-1000B) to extract system-agnostic insights. These insights reveal underlying patterns behind seemingly random data movement, providing actionable guidance for designing diverse future MoE serving systems. We validate our insights on a wafer-scale GPU architecture and achieve significant performance improvements with minimal modifications.

### v4

@@
-Orders in Chaos: Enhancing Large-Scale MoE LLM Serving with Data Movement Forecasting
+
 Abstract
@@
 To understand the patterns underlying this data movement, we conduct comprehensive data-movement-centric profiling across four state-of-the-art large-scale MoE models released in 2025 (200B-1000B) using over 24,000 requests spanning diverse workloads.
-We perform systematic analysis from both temporal and spatial perspectives and distill six key insights to guide the design of diverse future serving systems.
-With our insights, we then demonstrate how to improve wafer-scale GPUs as a case study, and show that minor architectural modifications leveraging the insights achieve substantial performance gains, delivering 5.3× and 3.1× average speedups on DeepSeek V3 and Qwen3 respectively.
-Our work presents the first comprehensive data-centric analysis of large-scale MoE models and a concrete design study using the learned lessons, with profiling traces and simulation framework already open-sourced with 1k downloads.
-Our traces and results are publicly available at https://huggingface.co/datasets/core12345/MoE_expert_selection_trace.
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+We perform systematic analysis from both temporal and spatial perspectives and distill six key insights to guide the design of diverse serving systems.
+We verify these insights on both future wafer-scale GPU architectures and existing GPU systems.
+On wafer-scale GPUs, lightweight architectural modifications guided by our insights yield a 6.6 average speedup across four 200B–1000B models.
+On existing GPU systems, our insights drive the design of a prefill-aware expert placement algorithm that achieves up to 1.25 speedup on MoE computation.
+Our work presents the first comprehensive data-centric analysis of large-scale MoE models together with a concrete design study applying the learned lessons.
+Our profiling traces are publicly available at https://huggingface.co/datasets/core12345/MoE_expert_selection_trace.
 I Introduction
-Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse domains, including programming assistance [llm_programming1, llm_programming2], translation [gpt_llm_trans1, lu2024llm_trans2], and chatbots [llm_chatbot1, llm_chatbot2].
-Since the beginning of 2025, large-scale Mixture of Experts (MoE) LLMs (200B+ model with 100+ experts) have become the leading models for frontier LLMs [chiang2024chatbot] and the most widely used open weight models.
+Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse domains, including programming assistance [47, 76], translation [1, 41], and chatbots [11, 71].
+Since the beginning of 2025, large-scale Mixture of Experts (MoE) LLMs (200B+ model with 100+ experts) have become the leading models for frontier LLMs [7] and the most widely used open weight models.
 Unlike dense LLMs that activate all model weights uniformly, MoE models dynamically route each token to only a subset of experts, introducing substantial data movement overhead.
-Such overhead already exceeds 50% of execution time for small models (e.g., Mixtral 8x7B) on modest systems (2-4 GPUs), and it exacerbates further with larger models, such as DeepSeek V3 with 32 experts and 15 parameters deployed on multi-node systems (32+ GPUs) [go2025moetuner, zhang2025comet].
+Such overhead already exceeds 50% of execution time for small models (e.g., Mixtral 8x7B) on modest systems (2-4 GPUs), and it exacerbates further with larger models, such as DeepSeek V3 with 32 experts and 15 parameters deployed on multi-node systems (32+ GPUs) [20, 89].
 However, as shown in Figure 1, no prior work has systematically investigated the underlying data movement patterns in large-scale MoE models.
-Prior studies conducted before the emergence of large-scale MoEs [fang2025klotski, tairin2025emoe, yao2024exploiting] have confined themselves to profiling much smaller models on limited hardware, typically examining only one or two LLMs while reporting simple observations without system-level insights.
+Prior studies conducted before the emergence of large-scale MoEs [15, 64, 84] have confined themselves to profiling much smaller models on limited hardware, typically examining only one or two LLMs while reporting simple observations without system-level insights.
 As parameter size and expert count surge, new patterns have emerged but remain unexplored, leaving significant optimization opportunities on the table.
@@
 To uncover the inherent patterns in MoE models, we conduct a comprehensive data-movement-centric profiling of four state-of-the-art MoE models ranging from 235B to 1000B parameters released in 2025.
-As highlighted in Figure 1, we profile DeepSeek V3 [liu2024deepseek], Llama4 Maverick [llama4], Qwen3-235B [yang2025qwen3], and Kimi K2 [team2025kimi] across 24,000 requests involving varied tasks, topics, and languages, which consumes 2000 GPU hours in total.
+As highlighted in Figure 1, we profile DeepSeek V3 [40], Llama4 Maverick [45], Qwen3-235B [82], and Kimi K2 [69] across 24,000 requests involving varied tasks, topics, and languages, which consumes 2000 GPU hours in total.
 We then collect the expert selection trace of all layers and tokens in each request to create an expert selection database of over 150 GB JSON files.
@@
 Our work represents the first systematic effort to characterize data movement patterns at the scale of up-to 1000B model across a wide range of tasks, providing actionable insights that can guide the design of next-generation MoE serving systems.
-To demonstrate the broad applicability of our insights, we present a case study on future GPU architecture design.
-Modern GPUs have adopted multi-chiplet designs due to single-die size limitations [mi300, cpelide, blackwell] and are evolving toward wafer-scale architectures enabled by emerging on-wafer packaging technologies [he2025waferlm, xu2025wsc].
-By applying our insights, we develop two key innovations for future wafer-scale GPUs: (1) A task distribution algorithm with a two-level data-placement-aware command processor architecture that balances workload across dies, and (2) a data-driven predictor with hardware-managed HBM architecture that reduces inter-die communication overhead by automatically caching critical data locally.
-These techniques achieve an average 4.0-6.3 speedup in MoE LLM serving throughput on wafer-scale GPUs.
+To demonstrate the broad applicability of our insights, we present case studies on both future and existing GPU systems.
+On the architecture side, we observe that modern GPUs have already adopted multi-chiplet designs due to single-die size limitations [60, 10, 48] and are evolving toward wafer-scale integration enabled by emerging on-wafer packaging technologies [23, 81].
+Targeting this trend, we develop a two-level data-placement-aware command processor and a hardware-managed HBM scheme that jointly balance workload across dies and reduce inter-die communication, achieving an average 6.6 speedup in MoE serving throughput on wafer-scale GPUs.
+On existing multi-GPU systems, we observe that prefill-stage expert selections can effectively predict decode-stage behavior.
+Building on this observation, we propose prefill-aware expert placement algorithms to reduce decode workload imbalance, and achieve up to 1.25 speedup.
 Our main contributions can be summarized as follows:
@@
 We distill six key insights for designing efficient MoE serving systems based on our profiling and analysis, providing actionable guidance that can inspire future research in MoE serving systems.
-Leveraging these insights, we conduct a case study on wafer-scale GPU architecture design. We proposed two strategies that improve MoE serving performance by 4.0-6.3 with minor hardware modifications.
+Leveraging these insights, we conduct case studies on both future and existing GPU systems.
+On future wafer-scale GPUs, we improve MoE throughput by 6.6 with minor hardware modifications.
+On existing multi-GPU systems, we achieve up to 1.25 speedup on an 8H100.
 We collect over 70,000 expert selection traces across multiple models and datasets, totaling over 150 GB in JSON format, and have open-sourced (1k downloads) all traces with our multi-chiplet simulator to facilitate future research on MoE serving systems.
@@
 II-A LLM and MoE Model Architecture
-Most state-of-the-art LLMs adopt a decoder-only transformer architecture that follows a token-by-token auto-regressive workflow [attention].
+Most state-of-the-art LLMs adopt a decoder-only transformer architecture that follows a token-by-token auto-regressive workflow [72].
 As shown in Figure 3(a), after users input queries, the serving process is divided into two stages: the prefill stage and the decode stage.
@@
 The generated token from each iteration is appended to the input sequence to produce the next token in the following iteration.
-The Mixture of Experts (MoE) mechanism is a state-of-the-art approach to improve LLM performance and has become prevalent among current frontier LLMs [moe-survey].
+The Mixture of Experts (MoE) mechanism is a state-of-the-art approach to improve LLM performance and has become prevalent among current frontier LLMs [44].
 As shown in Figure 3(b), MoE-based LLMs replace the feed-forward network (FFN) layers in traditional LLMs with MoE layers.
@@
 To address this overhead, existing research on MoE serving optimization has developed numerous system-level solutions targeting different performance and cost objectives.
-Edge systems like MoE-Lightning [moelightning] and CoServe [suo2025coserve] employ CPU memory offloading techniques to address GPU memory capacity constraints, while server systems such as Comet [zhang2025comet] and MegaScale-Infer [zhu2025megascale-infer] leverage multi-GPU architectures and address GPU-GPU communications in MoE for higher throughput.
-Novel hardware architectures like Duplex [yun2024duplex] explore processing-in-memory to accelerate data movement in MoE LLMs.
+Edge systems like MoE-Lightning [5] and CoServe [63] employ CPU memory offloading techniques to address GPU memory capacity constraints, while server systems such as Comet [89] and MegaScale-Infer [92] leverage multi-GPU architectures and address GPU-GPU communications in MoE for higher throughput.
+Novel hardware architectures like Duplex [87] explore processing-in-memory to accelerate data movement in MoE LLMs.
 However, these prior studies employ a system-centric methodology when optimizing for MoE LLMs.
@@
 III-C1.
-To quantify these relationships, we analyze expert pair frequencies across multiple requests and all model layers.
-The cumulative percentage of the most frequently selected expert pairs is shown in Figure 4(c).
-The results demonstrate that the top 20% of expert pairs account for 45%, 68%, 80%, and 55% of all activations in Deepseek, Qwen, Llama4, and Kimi K2, respectively.
+To quantify these relationships, we analyze expert pair frequencies across requests and layers.
+Figure Figure 4(c) shows that the top 20% of expert pairs account for 45%, 68%, 80%, and 55% of activations in Deepseek, Qwen, Llama4, and Kimi K2, respectively.
 This reveals strong, model-dependent cross-layer correlations, with Llama4111Llama4 places dense FFN layers between MoE layers, so we analyze correlations between adjacent MoE layers ( and ).
-exhibiting the strongest effect and Deepseek the weakest.
+showing the strongest effect and Deepseek the weakest.
 III-B2 Token-Level
@@
 Building on the layer-level and token-level relations, we observe notable similarities in expert selection patterns between prefill and decode stages.
-Comparing heatmaps across stages in Figure 6(a)(b)(c)(d), we find similar distributions of bright dots, indicating that expert pairs frequently selected during prefill are also likely selected during decode.
-This cross-stage consistency suggests expert selection in the decode stage can be predicted using prefill data.
+Comparing heatmaps across stages in Figure 6(a)(b)(c)(d), we find similar distributions of bright dots, indicating that expert pair heatmap during prefill and decode shares similarities.
+This cross-stage consistency suggests us that the prefill-collected information can guide initial decode steps until sufficient decode data accumulates.
 To quantify this similarity, we compute Spearman’s ratio () across all model layers, comparing prefill and decode heatmaps.
 Spearman’s Ratio measures monotonic relationships between variables, ranging from (perfect negative correlation) to (perfect positive correlation).
-Generally, indicates strong correlation, indicates moderate correlation, and suggests weak correlation [spearman].
+Generally, indicates strong correlation, indicates moderate correlation, and suggests weak correlation [53].
 The results in Figure 6(e)(f) show that most layers demonstrate strong correlation, while a few show moderate correlation.
 This makes it possible to predict decode-stage expert selection with prefill-stage data.
+Beyond expert-pair heatmaps, we also identify prefill-to-decode correlation at the single-expert frequency level.
+As shown in Figure 7(a), the frequency distributions of prefill and decode stages are substantially similar, though some discrepancies exist among low-frequency experts.
+To examine the most popular experts, we report the overlap rate of top experts between stages in Figure 7(b): the top-5 prefill experts cover around 60% of the top-5 decode experts, rising to 75% and 90% for top-10 and top-20, respectively.
+This indicates that prefill information can help predict the hottest decode experts.
+The cross-model Spearman correlation in Figure 7(c) confirms this relationship holds across all four models.
 III-B4 System Insights from Temporal Relation
@@
 This workload imbalance suggests system designs should duplicate or decentralize frequently used experts.
-To investigate selection patterns across different tasks, we analyze all 57 MMLU subjects spanning diverse fields, including biology, history, and math, etc [mmlu].
+To investigate selection patterns across different tasks, we analyze all 57 MMLU subjects spanning diverse fields, including biology, history, and math, etc [25].
 Figure 8(b) shows the top 10 most popular experts for each subject.
 Horizontal bright lines indicate certain experts are consistently activated regardless of subject, while remaining popular experts vary significantly between subjects, demonstrating both overlap and distinction in task-based expert selection.
-We further examine task impact using Chinese version of MMLU in MMLU Pro [mmlu_pro] with identical questions but different languages.
+We further examine task impact using Chinese version of MMLU in MMLU Pro [75] with identical questions but different languages.
 Figure 8(c) reveals distinctly different patterns: although 5-6 experts remain popular across subjects, only two overlap with English MMLU’s most frequently selected experts.
@@
 We follow Insight 3 to design a task allocation algorithm and leverage the temporal relation insights (Insight 1 and Insight 2) to build a data-driven predictor.
-We also make slight architectural modifications to support the task allocation algorithm and the data-driven predictor.
+We also make slight architectural modifications to support the proposed strategies.
 IV-A Trend of Future GPU Architecture
 GPU vendors are increasingly adopting multi-chiplet architectures to overcome single-die performance limitations.
-As Moore’s Law approaches its limits [lundstrom2003moore] and single-die size remains constrained by photomask dimensions (800-1,000 mm2), advanced packaging technologies like TSMC’s CoWoS [hu2023cowos], Samsung’s X-Cube [xcube], and Intel’s EMIB [emib] enable multiple chiplets within a single package.
-Leading vendors have adopted such designs: AMD’s MI300 [mi300] integrates eight compute chiplets, NVIDIA’s Blackwell features two chiplets [blackwell], and upcoming Rubin expects four [rubin].
-This trend is evolving toward wafer-scale systems [cowos2].
-TSMC’s System-on-Wafer (SoW) technology accommodates up to 24 compute dies and 96 HBM dies on a single wafer, exceeding 200,000 mm2 [sow-x].
+As Moore’s Law approaches its limits [42] and single-die size remains constrained by photomask dimensions (800-1,000 mm2), advanced packaging technologies like TSMC’s CoWoS [28], Samsung’s X-Cube [58], and Intel’s EMIB [43] enable multiple chiplets within a single package.
+Leading vendors have adopted such designs: AMD’s MI300 [60] integrates eight compute chiplets, NVIDIA’s Blackwell features two chiplets [48], and upcoming Rubin expects four [49].
+This trend is evolving toward wafer-scale systems [26].
+TSMC’s System-on-Wafer (SoW) technology accommodates up to 24 compute dies and 96 HBM dies on a single wafer, exceeding 200,000 mm2 [57].
 As shown in Figure 10(a), a typical wafer-scale multi-chiplet GPU consists of multiple units, each containing a GPU die and several HBM dies interconnected in a mesh topology.
 Such systems contain over 10 TB of HBM and PFLOPS-level computing power, supporting extremely large models and batch sizes.
-The TSMC SoW technology shown in Figure 9(b) connects each GPU die to local HBM dies through local-silicon interconnects (LSI) [lsi], with adjacent GPU dies communicating via LSI vertically or XSR SerDes links horizontally.
+The TSMC SoW technology shown in Figure 10(b) connects each GPU die to local HBM dies through local-silicon interconnects (LSI) [27], with adjacent GPU dies communicating via LSI vertically or XSR SerDes links horizontally.
 Although LSI and SerDes both provide terabit-level bandwidth, inter-GPU-die communication remains the primary bottleneck.
@@
 Simultaneous remote HBM access by multiple dies creates bandwidth contention and traffic congestion, further degrading performance.
-IV-B Challenges of serving MoE with future GPUs
+IV-B Background on Wafer-scale GPU Programming model
+The programming model of future wafer-scale chip remains an open question, but two major candidates have emerged: the multi-GPU-like and the single-GPU-like programming model.
+Multi-GPU-like programming model.
+WSC-LLM [81] and MoEntwine [68] adopt a multi-GPU-like programming model that exposes the entire wafer as a multi-GPU system.
+Programmers can program the wafer similarly to a conventional multi-GPU system, with the key difference being the 2D mesh topology where each die communicates directly only with its neighbors.
+While this approach offers fine-grained control over individual dies and enables flexible software strategies, it diverges from current industry trends.
+For instance, although Blackwell and Rubin both integrate two compute dies, NVIDIA exposes no toolchain for fine-grained die-level control.
+Multi-Instance GPU (MIG) can partition a multi-die GPU into independent GPU instances, making each die act independently, but the high-speed D2D link is disabled in this mode [34], forcing inter-die communication through the 10-100 slower NVLink or PCIe and negating the benefit of multi-die packaging.
+Therefore, extending this programming model to wafer-scale GPUs would require substantial architectural changes to current GPU designs, including redesigning the D2D/C2C controller workflow and distributed LLC structure, making it infeasible in the near term.
+Single-GPU-like programming model.
+HDPAT [80], Hecton [29], and our work adopt a single-GPU-like programming model that presents the entire wafer as a unified GPU, completely abstracting the multi-die topology and data placement from software.
+All HBM stacks form a unified memory space, and the GPU resolves data locations directly via physical addresses, making the programming experience identical to that of a monolithic GPU.
+We adopt this model for two key reasons.
+First, it aligns with commercial multi-chiplet GPUs such as Blackwell and Rubin, ensuring practical industry relevance.
+Second, it eliminates the programming complexity of explicit inter-die communication management through libraries such as NCCL or NVSHMEM.
+However, this programming simplicity transfers the optimization burden entirely to the hardware.
+Given the inherently distributed architecture, local versus remote data access costs vary by up to 15, yet the abstraction prevents programmers from controlling cross-die data movement.
+Consequently, architecture-level optimizations become critical to achieve high hardware utilization.
+IV-C Challenges of serving MoE with future GPUs
 Unlike current multi-GPU systems, wafer-scale GPUs can fit entire MoE models on a single chip and support batch sizes over 10,000.
@@
 Current GPUs integrate a CPU in their SoC to serve as a command processor and allocate tasks to all SMs.
-However, the traditional command processors treat all SMs equally, ignoring their physical locations and data placement [cp1, cp2].
+However, the traditional command processors treat all SMs equally, ignoring their physical locations and data placement [36, 22].
 This oblivious task-to-SM assignment generates excessive D2D traffic and ignores MoE expert selection skewness, leading to poor utilization when most dies remain idle while others become overloaded.
@@
 Frequently accessed experts in remote HBM could be cached locally to minimize D2D traffic, but current GPUs do not distinguish between local and remote HBM and therefore generate unnecessary traffic.
-IV-C Motivation and Insights
+IV-D Motivation and Insights
 To address these challenges, we propose two strategies with architectural support.
@@
 If future programming models provide more fine-grained features, these strategies could alternatively be implemented at the system level without architectural changes.
-IV-D Architecture Design
-IV-D1 Architecture Overview
+IV-E Architecture Design
+IV-E1 Architecture Overview
 To support our proposed predictor and task allocation algorithm, we implement two architectural modifications with minimal overhead, as illustrated in Figure 10(a).
@@
 These enable autonomous caching of remote data in local HBM and intelligent redirection of data requests, reducing inter-die communication overhead.
-IV-D2 Key Data Structures
+IV-E2 Key Data Structures
 There are two key data structures: Global CP data and the PDU prediction table.
@@
 The cross-token heatmap records expert activation patterns over time, providing historical data for prediction generation.
-PDU Prediction Table: Each PDU stores a prediction table with two key fields per expert: the cp_en bit indicating whether the expert should be cached locally (calculated by Global CP and transferred before computation), and the is_local bit tracking whether the expert is already cached (maintained by PDU itself).
-IV-D3 Workflow During Kernel Launch
-When a new kernel launches ( ), the Global CP implements our task allocation algorithm to split the MoE kernel into sub-kernels for each die and executes the predictor algorithm to generate duplication guidance (cp_en field is PDU).
+PDU Prediction Table: Each PDU stores a prediction table with two key fields per expert: the cp_en bit indicating whether the expert should be cached locally (calculated by Global CP and transferred to each die), and the is_local bit tracking whether the expert is already cached in local HBM.
+IV-E3 Workflow During Kernel Launch
+When a new kernel launches ( ), the Global CP runs our task allocation algorithm to split the MoE kernel into per-die sub-kernels and executes the predictor to generate duplication guidance (cp_en field is PDU).
 The Global CP then sends sub-kernels and prediction information to local CPs ( ).
-After receiving the guidance information from Global CP, each local CP allocates tasks to its SMs ( ) and configures the prediction table in the D2D controller for local HBM management ( ).
-After computation, the local CP collects expert duplication statistics and sends them to the Global CP to update expert distribution information.
-IV-D4 Dataflow for Remote Data Access
+Each local CP allocates tasks to its SMs ( ) and configures the prediction table in the D2D controller for local HBM management ( ).
+After computation, local CPs collect expert duplication statistics and send them to Global CP to update expert distribution information.
+IV-E4 Dataflow for Remote Data Access
 We integrate ATU and PDU into the D2D controller to support hardware-managed HBM by modifying the remote data access flow.
@@
 The LLC and memory controller then retrieve data and send it back to the SM( , ).
-IV-D5 Algorithm Design
+IV-E5 Algorithm Design
 This subsection presents our task allocation algorithm and data-driven predictor, both implemented by the global CP.
@@
 V-A Experiment Setup
-Methodology: We conduct experiments using trace-driven simulation on a validated simulator.
-Expert selection traces are collected by deploying SGLang [zheng2024sglang] on an 8×H100 DGX server and an 8×H200 AWS instance.
+Methodology: We conduct experiments using event-driven simulation on a validated simulator.
+Expert selection traces are collected by deploying SGLang [90] on an 8×H100 DGX server and an 8×H200 AWS instance.
 We developed a custom multi-chiplet GPU simulator in Python since existing tools are inadequate for our needs.
-Cycle-accurate simulators like Gem5 [gem5], gpgpusim [gpgpusim], and mgpusim [mgpusim] accurately model single GPUs but are too slow for large-scale systems with 20+ GPUs serving batch sizes exceeding 15,000.
-Event-driven ASTRA-sim [astra] supports multi-GPU systems but lacks MoE serving support and detailed single-GPU microarchitecture modeling.
+Cycle-accurate simulators like Gem5 [4], gpgpusim [2], and mgpusim [62] accurately model single GPUs but are too slow for large-scale systems with 20+ GPUs serving batch sizes exceeding 15,000.
+Event-driven ASTRA-sim [78] supports multi-GPU systems but lacks MoE serving support and detailed single-GPU microarchitecture modeling.
 To ensure accuracy, we validated our simulator against real data from an 8×H100 DGX server, as shown in Sec.
@@
 DRAM |
-| Dojo | 5 | 5 | 2 TB/s | 1.5 TB/s | 80GB | 1000 TFLOPS | ||||||
-| TSMC-SoW | 3 | 8 | 2 TB/s | 1.5 TB/s | 80GB | 1000 TFLOPS | ||||||
+| Dojo | 5 | 5 | 3.35 TB/s | 1.7 TB/s | 80GB | 989 TFLOPS | ||||||
+| TSMC-SoW | 3 | 8 | 3.35 TB/s | 1.7 TB/s | 80GB | 989 TFLOPS | ||||||
 | Dojo-Enhanced | 5 | 5 | 8 TB/s | 2 TB/s | 180GB | 4500 TFLOPS | ||||||
@@
 Metric: We measure pure MoE layer performance during the decode stage as modern LLM serving systems show a trend toward fine-granularity disaggregation.
-Traditional LLMs benefit from separating prefill and decode stages across different machines, as demonstrated by DistServe [zhong2024distserve] and subsequent works [loongserve, qin2025mooncake].
+Traditional LLMs benefit from separating prefill and decode stages across different machines, as demonstrated by DistServe [91] and subsequent works [79, 51].
 For MoE models, this disaggregation extends further.
-MegaScale-Infer [zhu2025megascale-infer] separates attention and MoE operations onto different machines for optimal batch sizes.
+MegaScale-Infer [92] separates attention and MoE operations onto different machines for optimal batch sizes.
 Following this trend, we focus on optimizing MoE operations during the decode stage.
-Hardware Configuration: We evaluate two multi-chiplet topologies: Tesla Dojo [dojo1, dojo2] and the TSMC SoW roadmap [tsmc-arch].
+Hardware Configuration: We evaluate two multi-chiplet topologies: Tesla Dojo [65, 66] and the TSMC SoW roadmap [70].
 As summarized in Table I, Dojo uses a 55 2D mesh, while TSMC SoW adopts an 83 2D mesh.
 These choices reflect a deployed system (Dojo) and near-future industry support (TSMC SoW).
-For both the Dojo and TSMC SoW configurations, each chiplet is A100-like, providing 1,000 TFLOPS FP8 compute, 80GB HBM, 2TB/s local HBM bandwidth, and 1.5 TB/s inter-die bandwidth to adjacent chiplets.
+For both the Dojo and TSMC SoW configurations, each chiplet is H100-like, providing 1,000 TFLOPS FP16 compute, 80GB HBM, 3.35TB/s local HBM bandwidth, and 1.7 TB/s inter-die bandwidth to adjacent chiplets.
 We also include an extended experiment in Sec.
@@
 Baseline Configurations: We compare our approach against the simple strategy currently used by GPU.
-The Base configuration assigns equal numbers of experts to each die but ignores expert placement when allocating tasks.
+The Base configuration adopts an EP-like data placement and assigns an equal number of experts to each die.
+However, the entire wafer operates as a single large GPU: each die handles the same amount of expert computation without considering expert placement.
+EP assigns each expert’s computation to the die where it resides, as also adopted by MoEntwine [68].
+This eliminates all D2D communication but can cause severe workload imbalance.
+Note that even under EP, our Global CP and Local CP architecture remains necessary, as expert placement information is still required.
 We implement three variants: Allo Only uses solely our task allocation strategy; Pred Only includes only the data-driven predictor; and Allo+Pred combines both techniques.
@@
 Models and Workloads: We conduct evaluations with real traces collected from Qwen3 and Deepseek V3.
-The traces are gathered from diverse datasets, including MMLU [mmlu], MMLU Pro [mmlu_pro], ChineseSimpleQA [he2024chinese], and LiveCodeBench [jain2024livecodebench], comprising over 24,000 requests per model.
+The traces are gathered from diverse datasets, including MMLU [25], MMLU Pro [75], ChineseSimpleQA [24], and LiveCodeBench [32], comprising over 24,000 requests per model.
+Each test batch is filled by sequentially adding requests in the order of MMLU, MMLU-Pro (CH), ChineseSimpleQA, and LiveCodeBench until the target batch size is reached.
 V-B Validation of Simulator
@@
 We evaluate both single-GPU execution and two-GPU peer-to-peer (P2P) communication.
-For single-GPU execution, we benchmark GEMM operations for experts in the MoE layer across varying batch sizes for both DeepSeek and Qwen.
+For single-GPU execution, we benchmark one expert in a MoE layer, which consists of three GEMM operations, across varying batch sizes for both DeepSeek and Qwen.
 For P2P communication, we measure data migration between two GPUs over payload sizes ranging from 4 KB to 4 GB.
@@
 We evaluate MoE decode stage throughput in Figure 12, with results normalized to the baseline configuration.
-Comparison across models: Our Allo+Pred strategy achieves 5.3 throughput improvement on Deepseek and 3.1 on Qwen.
-Deepseek shows greater improvement due to its larger number of experts (256 vs.
-128) and more complex expert selection patterns.
-Without our expert placement-aware allocation, serving Deepseek introduces significantly more inter-unit communication than Qwen.
-Since our task allocation strategy places computation close to data, it yields more substantial improvements on Deepseek.
-Comparison across chiplet architectures: Our Allo+Pred strategy shows 3.7 improvement on Dojo and 4.7 on TSMC, despite similar die counts (25 vs.
+Comparison across models: Our Allo+Pred strategy achieves 7.0, 8.2, 7.3, and 4.1 throughput improvement on Deepseek, Kimi, Llama, and Qwen, respectively.
+Deepseek and Kimi benefit more due to their larger expert count (256 vs.
+128) and more complex selection patterns.
+Comparison across chiplet architectures: Our strategy shows 6.0 improvement on Dojo and 7.5 on TSMC, despite similar die counts (25 vs.
 24).
-The improvement on TSMC is more substantial due to its rectangular layout, where dies are farther apart, introducing more inter-unit communication without strategic task allocation.
+TSMC’s rectangular layout places dies farther apart, introducing more inter-unit communication without strategic task allocation, hence the larger gain under our strategy.
+Comparison with EP: At small batch sizes such as 4096, our strategy and EP perform similarly: few tokens per expert make execution memory-bound, so splitting one expert across multiple dies offers no benefit, and our strategy degenerates to EP.
+The advantage emerges at larger batches, achieving 1.44 speedup over EP at batch size 16,384.
 V-D Hop Reduction
@@
 Higher hop counts indicate frequent cross-die data movement.
-We normalize results to baseline and report hop reduction ratios, where a ratio of 10 means hop count is reduced to 1/10.
-Pred Only reduces hop counts by 3.7-4.4, aligning with performance results in Figure 11.
+We normalize results to baseline and report hop reduction ratios, where a ratio of 10 means the hop count is reduced to 1/10.
+Pred Only reduces hop counts by 4.5, aligning with performance improvement of 3.0.
 This indicates cross-unit communication is the primary bottleneck in baseline, and reducing hop counts proportionally improves performance.
-Allo Only reduces hop counts by 7.2-10.8, exceeding the performance improvement range of 17-661.
+Allo Only reduces hop counts by 142, exceeding the performance improvement of 6.3.
 This shows that with our allocation algorithm, inter-unit communication is no longer the sole bottleneck.
 While reducing hop counts still improves performance, the improvement is not proportional.
-Allo+Pred reduces hop counts by over 210-944 compared to baseline and achieves 10 improvement over Allo Only.
-However, performance improvement is only 2.5-6.5 over baseline, with just 1.2 average improvement over Allo Only.
+Allo+Pred reduces hop counts by over 213 compared to baseline.
+However, performance improvement is only 6.63 over baseline, with just 1.1 average improvement over Allo Only.
 This demonstrates that hop count is no longer a performance bottleneck.
-Our task allocation successfully places computation near relevant data.
-Most tasks are distributed to local dies holding related experts, with only extremely popular experts requiring remote allocation.
+With the help of our task allocation algorithm, most tasks are distributed to local dies holding related experts, with only extremely popular experts requiring remote allocation.
 This leads to minimal D2D traffic and shifts the bottleneck to workload distribution.
@@
 As GPU performance outpaces interconnect bandwidth, implementing the allocator in the GPU command processor becomes increasingly necessary to sustain performance.
-V-G Discussion
-This wafer-scale GPU architecture serves as a case study demonstrating the practical applicability of our proposed insights, which constitute the paper’s primary contribution.
-Our design implements only a subset of the proposed insights to validate their effectiveness.
-We follow Insight 3 for task allocation and leverage temporal relation insights (Insight 1 and part of Insight 2) to build a data-driven predictor.
-However, the remaining insights offer significant potential for further exploitation in wafer-scale architectures.
-For instance, Insight 2 could inform LLC prefetch strategies and enhance task allocation algorithms that consider LLC data placement.
-Spatial relation insights (Insight 4 and Insight 5) could optimize initial expert placement, which could be further refined by incorporating task information from Insight 6.
-Importantly, our insights extend far beyond this case study to benefit diverse future systems, including multi-GPU clusters [zhu2025megascale-infer, zhang2025comet], CXL-based or CPU-based memory disaggregation [moelightning, fang2025klotski], Flash-based multi-hierarchy systems [cambricon-llm, sheng2023flexgen], and other emerging architectures and systems.
-VI Related Work
+V-G Area and Power Overhead
+We estimate the area and power overhead of all added modules in Table II.
+Our parameters support up to 100 layers with 512 experts per layer, well beyond the largest existing MoE model (Kimi-K2: 61 layers, 384 experts).
+The full heatmap for 100 layers totals 50MB and is stored in Global CP DRAM, with an on-chip heatmap cache of 0.5MB (512-bit width) buffering one layer at a time.
+The Prediction Table is implemented in registers due to its small size, while all other components use SRAM.
+Register synthesis is performed with Yosys [77], and SRAM is modeled with CACTI [3], both scaled to 5nm to match the H100 process node.
+Area estimates for the Global CP and Local CPs are derived from ARM core data.
+As shown in Table II, the total area and power overhead is less than 0.04%.
+| Capacity |
+| Prediction Table | 128 B | 16 bit | 25 | 0.0020 | 55.75 | ||||||||
+| Address Translation Unit | 4.25 KB | 68 bit | 25 | 0.0048 | 334.25 | ||||||||
+| Local CP (A72) [54] | N/A | N/A | 25 | 7.5000 | 7000 | ||||||||
+| Expert Distribution Table | 4.5 KB | 72 bit | 1 | 0.0002 | 13.94 | ||||||||
+| Heatmap Cache | 0.5 MB | 512 bit | 1 | 0.0278 | 184.67 | ||||||||
+| Global CP (A76) [74] | N/A | N/A | 1 | 1.1000 | 1000 | ||||||||
+| Total | 6.13 | 8588.61 | |||||||||||
+| Overhead (25-die wafer) | 0.04% | 0.04% |
+V-H Discussion
+insights, which constitute the paper’s primary contribution.
+Specifically, we follow Insight 3 for task allocation and leverage temporal relation insights (Insight 1 and part of Insight 2) to build a data-driven predictor.
+However, the remaining insights offer significant potential to further enhance the performance of wafer-scale GPUs in future works.
+For instance, Insight 2 could inform LLC prefetch strategies and enhance task allocation with LLC data placement awareness; spatial relation insights (Insight 4 and Insight 5) could optimize initial expert placement, further refined by task information from Insight 6.
+Importantly, our insights extend far beyond this case study to benefit diverse future systems, including multi-GPU clusters [92, 89], CXL-based or CPU-based memory disaggregation [5, 15], Flash-based multi-hierarchy systems [86, 56], and other emerging architectures.
+VI Case Study2: Decode Expert Placement with Prefill Information on Current GPUs
+VI-A Introduction
+Workload imbalance is one of the biggest challenges in large-scale MoE serving (200+ GPUs).
+EPLB [12] addresses this by dynamically adjusting expert placement, but it triggers every 3000+ steps and relies on periodically collected profiling data [73].
+A natural question then arises: how to set expert placement for the initial 1000 decode tokens when no profiling data are yet available?
+This is especially pressing for short-output requests, for which EPLB never collects enough data to be effective.
+Inspired by Insight 1, which reveals temporal correlation between prefill and decode stages, we propose leveraging prefill-stage expert selection information to guide expert placement for initial decode steps.
+As shown in Figure 16, we design two placement algorithms (details in algorithm 2).
+The Remap-based algorithm keeps the number of experts per GPU unchanged and reassigns experts across GPUs for a more balanced workload: it sorts experts by decreasing roofline cost and greedily assigns each to the least-loaded GPU, subject to a uniform capacity of experts per GPU.
+The Duplication-based algorithm reserves extra expert slots on each GPU and uses prefill traces to duplicate hot experts, thereby avoiding congestion: starting from the default contiguous layout (e.g., experts 0–15 on GPU,0, 16–31 on GPU,1, etc.), it greedily adds up to extra replicas per GPU, selecting at each step the (expert, GPU) pair that maximally reduces the bottleneck load ; tokens of a replicated expert are evenly split among all its copies.
+Both algorithms use a roofline-based cost model to estimate per-GPU load.
+VI-B Methodology
+We deploy Qwen3-235B with SGLang on 8H100 GPUs with NVLink.
+We build a distributed profiler by inserting cuda.Event timers into SGLang to measure individual operations (attention, top-, all-to-all, and MoE) on each GPU independently.
+We manipulate expert placement through SGLang’s init_expert_location interface and use DeepEP as the MoE backend.
+The ep_dispatch_algorithm is set to “dynamic” so that tokens are evenly distributed across replicas of a duplicated expert.
+Metric. We report MoE computation time, i.e., all three expert linear layers, excluding attention, all-to-all, and top-.
+Model and Benchmark.
+We evaluate on Qwen3-235B (94 MoE layers, 128 experts per layer, 8 selected).
+We use MMLU and Global-MMLU datasets, following the original ordering.
+Batch sizes range from 64 to 16,384.
+Baselines.
+Default is the standard contiguous placement used by Qwen and SGLang (experts 0–15 on GPU-0, 16–31 on GPU-1, etc.).
+Best and Worst are the theoretically optimal and worst placements generated with oracle decode-stage selections (not available in practice).
+Remap and Dup are our two prefill-guided strategies.
+For Dup, we use one extra slot per GPU, yielding 128+8=136 experts per layer.
+VI-C Results
+As shown in Figure 17, Remap and Dup achieve speedups of 15.5% and 12.5% over Default, respectively, and deliver over 2 speedup compared with Worst.
+Both remain within 10% of Best, which exploits decode-stage information unavailable in practice, demonstrating the effectiveness of our approach.
+Since the two algorithms perform comparably, one can choose between them to fit different memory and system constraints.
+We note that our 8-GPU EP scale inherently limits the achievable improvement: with EP8, each GPU holds 16 experts per layer, so every GPU likely contains a mix of hot and cold experts, naturally yielding a relatively balanced workload even under the Default layout (the max/min execution-time ratio is only about 1.3).
+We expect greater speedups at larger EP scales where load imbalance is more pronounced.
+VII Related Work
 MoE model behavior studies:
-Several MoE model tech reports [muennighoff2024olmoe, jiang2024mixtral, dai2024deepseekmoe, sun2024hunyuan] provide the MoE routing patterns as part of their evaluation.
-For example, the Mixtral report [jiang2024mixtral] shows the temporal locality of expert selection by reporting the percentage of repetitive assignment.
-The OLMoE report [muennighoff2024olmoe] shows the co-activation pattern and domain specialization among the experts.
-A blog post from SGLang [sglang-deepseek-blogpost] shows expert distribution statistics for the DeepSeekV3 model and the inherent imbalance in expert selection and similarity between prefill and decode.
+Several MoE model tech reports [46, 33, 9, 61] provide the MoE routing patterns as part of their evaluation.
+For example, the Mixtral report [33] shows the temporal locality of expert selection by reporting the percentage of repetitive assignment.
+The OLMoE report [46] shows the co-activation pattern and domain specialization among the experts.
+A blog post from SGLang [18] shows expert distribution statistics for the DeepSeekV3 model and the inherent imbalance in expert selection and similarity between prefill and decode.
 None of these studies provides a comprehensive profiling across multiple large-scale (200B) MoE models, nor do they present a data-movement-centric methodology to highlight the opportunities like this paper.
 Data movement optimization for MoE LLM inference
-Various prior work [hwang2024pre, li2023accelerating, zhai2023smartmoe, gupta2024lynx, kamahori2025fiddler, eliseev2023fast, du2024sida, li2024optimizing, skliar2024mixture, moelightning, yun2024duplex, zhang2025comet, hwang2023tutel] have focused on improving the efficiency of MoE LLM inference by reducing data movement.
-For example, Lina [li2023accelerating] exploits the skewness in expert selection to dynamically schedule resources during inference and balance traffic across multiple GPUs.
-LYNX [gupta2024lynx] dynamically reduces active experts while preserving model accuracy.
-Pre-gate MoE [hwang2024pre] uses a pre-gating function to alleviate the dynamic nature of expert selection.
-Sida [du2024sida] builds an offline hash function to predict expert usage to reduce data movement between CPU and GPU.
-MoE-Lightning [moelightning] leverages a CPU-GPU pipeline and paged weights to improve resource utilization.
-Eliseev and Mazur [eliseev2023fast] exploit expert locality and leverage LRU caching to manage GPU and CPU memory.
-This work aims to reduce data movement in MoE LLM inference of today’s systems, and our case study shows that data-driven profiling across multiple models provides a solid base to optimize not only current but future systems of any scale.
+Various prior work [31, 38, 21, 35, 14, 13, 37, 59, 5, 87, 89, 30] have focused on improving the efficiency of MoE LLM inference by reducing data movement.
+For example, Lina and SmartMoE [38, 88] exploit expert selection skewness to dynamically schedule resources during inference and balance traffic across GPUs.
+LYNX [21] dynamically reduces active experts while preserving model accuracy.
+Pre-gate MoE [31] uses a pre-gating function to alleviate the dynamic nature of expert selection.
+Sida [13] builds an offline hash function to predict expert usage and reduce data movement between CPU and GPU.
+MoE-Lightning [5] leverages a CPU-GPU pipeline and paged weights to improve resource utilization.
+Eliseev and Mazur [14] exploit expert locality and leverage LRU caching to manage GPU and CPU memory.
+This work targets data movement reduction in MoE LLM inference. Our cross-model profiling reveals optimization principles that apply broadly to current and future systems regardless of scale.
 Wafer-Scale and Chiplet Architectures
-As single chip scaling slows down, packaging technologies such as wafer-scale and chiplet packaging have become a promising path to further improve compute efficiency for future systems.
-Recent work in this area either focuses on the interconnect design [rashidi2025fred, yang2025pd, yu2025cramming, li2024lucie, chen2024waferscale, feng2023heterogeneous, feng2023scalable] of wafer-scale and chiplet systems, or
-general data placement for specific algorithms [he2025waferlm, tan2021nn] and applications [shao2019simba, xu2025wsc, odema2024scar].
-In contrast, our work is the first to study the implications of serving large-scale MoE LLMs in a wafer-scale GPU and propose data-movement-centric HW/SW co-design optimizations.
-VII Conclusion
-In this work, unlike previous MoE serving works that adopt system-centric approaches and develop deployment-specific strategies, we study MoE model serving from a model-focused perspective We conducted comprehensive data-movement-centric profiling of the latest large-scale MoE models (200B-1000B) to extract system-agnostic insights.
-These insights reveal underlying patterns behind seemingly random data movement, providing actionable guidance for designing diverse future MoE serving systems.
-We validate our insights on a wafer-scale GPU architecture and achieve significant performance improvements with minimal modifications.
+As single-chip scaling slows, wafer-scale and chiplet packaging offer promising paths for improved compute efficiency.
+Prior work targets either interconnect design [52, 83, 85, 39, 6, 17, 16] or data placement for specific algorithms [23, 67] and applications [55, 81, 50].
+In contrast, we are the first to study MoE LLM serving on wafer-scale GPUs and propose data-movement-centric HW/SW co-design optimizations.
+VIII Conclusion
+In this work, unlike prior MoE serving studies that adopt system-centric approaches and develop deployment-specific strategies, we study MoE serving from a model-focused perspective.
+We conduct comprehensive data-movement-centric profiling of the latest large-scale MoE models (200B–1000B) to extract system-agnostic insights.
+These insights reveal structured patterns underlying seemingly random data movement, providing actionable guidance for designing diverse future MoE serving systems.
+We validate our findings on both a future wafer-scale GPU architecture and existing multi-GPU systems, achieving significant performance improvements through minimal architectural modifications and lightweight software design, respectively, demonstrating the effectiveness and broad applicability of our insights.